OpenStudy (anonymous):

how to simplify radical 1/5

OpenStudy (unklerhaukus):

\[\sqrt{\frac 15}=\sqrt{\frac 15}\times 1=\sqrt{\frac 15}\times\sqrt\frac 5 5=\frac{\sqrt5}{5}\]

OpenStudy (anonymous):

OHHH LOL I'm thinking too complicated thanks!